### v1

Abstract
Let $G$ be a connected graph with order $n$ and size $m$. Let $D(G)$ and $Tr(G)$ be the distance matrix and diagonal matrix with vertex transmissions of $G$, respectively. For any real $\alpha\in[0,1]$, the generalized distance matrix $D_{\alpha}(G)$ of $G$ is defined as
$D_{\alpha}(G)=\alpha Tr(G)+(1-\alpha)D(G).$
The $D_{\alpha}$ spectral radius of $G$ is the spectral radius of $D_{\alpha}(G)$, denoted by $\mu_{\alpha}(G)$. In this paper, we establish a lower bound on the difference between the maximum vertex transmission and the $D_{\alpha}$ spectral radius of non-transmission regular graphs, and we also characterize the extremal graphs attaining the bound.
Key Words:  Generalized distance matrix, Spectral radius, Non-transmission regular graph, Lower bound. AMS Subject Classification (2020):  05C50,15A18
1 Introduction
Throughout this paper all graphs considered are simple and connected. Let $G=(V(G),E(G))$ be a graph with vertex set $,V(G)=\{v_{1},v_{2},$$\ldots,v_{n}\}$ and edge set $E(G)$. The number of vertices $|V(G)|=n$ is the order of $G$ and the number of edges $|E(G)|=m$ is the size of $G$. For any vertex $v_{i}\in V(G)$, $N(v_{i})=\{v_{j}:(v_{i},v_{j})\in E(G)\}$ is called the set of neighbors of $v_{i}$. The degree $d(v_{i})$ of a vertex $v_{i}$ in $G$ is the number of edges of $G$ incident with $v_{i}$, i.e. , $d(v_{i})=|N(v_{i})|$. Let $K_{n}$ denote the complete graph of order $n$. For other concepts, please refer to [4].
The distance between two vertices $v_{i}$ and $v_{j}$, denoted by $d(v_{i},v_{j})$ or simply by $d_{ij}$, is the length of a shortest path from $v_{i}$ to $v_{j}$ in $G$. The diameter of $G$, denoted by $diam(G)$, is the maximum distance between any two vertices of $G$. For any vertex $v_{i}\in V(G)$, the transmission of $v_{i}$ in $G$, denoted by $Tr_{i}$, is the sum of distances from $v_{i}$ to all other vertices of $G$, that is, $Tr_{i}=\sum\limits_{k=1}^{n}d_{ik}$. For a connected graph $G$, if $Tr_{1}=Tr_{2}=\cdots=Tr_{n}=r$, we say that $G$ is $r$-transmission regular; otherwise, $G$ is non-transmission regular. The second transmission degree of $v_{i}$, denoted by $T_{i}$, which is given by $T_{i}=\sum\limits_{t=1}^{n}d_{it}Tr_{t}$. The sum of all distances between all unordered pairs of vertices in $G$ is called the Wiener index of graph $G$, denoted by $W(G)$, i.e., $W(G)=\frac{1}{2}\sum\limits_{i=1}^{n}Tr_{i}$.
For a connected graph $G$, the distance matrix of $G$ denoted by $D(G)=(d_{ij})_{n\times n}$, where $d_{ij}=d(v_{i},v_{j})$ is the distance from vertex $v_{i}$ to vertex $v_{j}$. Let $Tr(G)=\textrm{diag}(Tr_{1},Tr_{2},\ldots,Tr_{n})$ be the diagonal matrix with vertex transmissions of $G$. With the development of the matrix theory of graphs, Aouchiche and Hansen [1] defined the distance signless Laplacian matrix of $G$ as $D^{Q}(G)=Tr(G)+D(G)$. The eigenvalues of $D(G)$ ($D^{Q}(G)$) are called the distance (distance signless Laplacian) eigenvalues of $G$, which are all real. The largest distance (distance signless Laplacian) eigenvalue of $G$ is called distance (distance signless Laplacian) spectral radius of $G$, denoted by $\mu(G)$ ($\mu_{Q}(G)$).
So far, the distance spectral radius and distance signless Laplacian spectral radius of a connected graph have been investigated extensively. For example, in [18], Nath and Paul determined the unique graph with minimum distance spectral radius among all connected bipartite graphs of order $n$ with a given matching number. In [12], Lin and Lu determined the graphs with minimum and maximum distance signless Laplacian spectral radius among connected graphs with fixed order and clique number. In [14], Lin and Zhou determined the unique trees with second and third maximum distance signless Laplacian spectral radius, respectively, and they also determined the unique connected graph with second maximum distance signless Laplacian spectral radius. In [24], Zhang et al. determined the minimum distance spectral radius of graphs among n-vertex graphs with given connectivity and independence number, and characterized the corresponding extremal graph. In [16], Liu et al. posed a conjecture about the difference between the maximum vertex transmission and the distance spectral radius. In [15], Liu et al. gave a lower bound for the difference between the maximum vertex transmission and the distance spectral radius. In [11], Lan gave a sharp lower bound of $2Tr_{max}-\mu_{Q}(G)$ among all $n$-vertex connected graphs, and characterized the extremal graphs. For more results about distance and distance signless Laplacian spectral radius of graphs, readers may refer to [2, 3, 9, 20, 21, 23] and the references therein.
In [6], Cui et al. proposed to study the convex linear combinations of $Tr(G)$ and $D(G)$, deonted by $D_{\alpha}(G)$ and defined as
$D_{\alpha}(G)=\alpha Tr(G)+(1-\alpha)D(G),\ \ \alpha\in[0,1].$
We call $D_{\alpha}(G)$ as the generalized distance matrix of $G$. It is easy to see $D_{0}(G)=D(G)$, $D_{1}(G)=Tr(G)$ and $2D_{\frac{1}{2}}(G)=D^{Q}(G)$, thus, the family $D_{\alpha}(G)$ extends both $D(G)$ and $D^{Q}(G)$. Therefore, it is very important to study the generalized distance matrix. The largest eigenvalue of $D_{\alpha}(G)$ is called the $D_{\alpha}$ spectral radius or generalized distance spectral radius of $G$, denoted by $\mu_{\alpha}(G)$.
Recently, the generalized distance spectral radius have been studied in some papers. In [7], Cui et al. characterized the unique graph with minimum generalized distance spectral radius among the connected graphs with fixed chromatic number. In [8], Diaz et al. presented some extremal results on the spectral radius of generalized distance matrix that generalize previous results on the spectral radii of the distance matrix and distance signless Laplacian matrix. In [13], Lin et al. gave several graph transformations on the $D_{\alpha}$-spectral radius. There are also some other interesting results on the $D_{\alpha}$ spectral radius which can be found in [5, 8, 17, 19].
In the theory of graph spectra, there has a classic problem which is establish upper or lower bounds for related spectral radius of given class of graphs. Motivated by [15] and [11], in this paper we decide to study the difference between the maximum vertex transmission and the $D_{\alpha}$ spectral radius of connected graphs, we establish a lower bound on the difference between the maximum vertex transmission and the $D_{\alpha}$ spectral radius of non-transmission regular graphs, and we also characterize the extremal graphs attaining the bound.
2 Lower bound on the difference between the maximum vertex transmission degree and the $D_{\alpha}$ spectral radius
In this section, we will establish a lower bound on the difference between the maximum vertex transmission and the $D_{\alpha}$ spectral radius for graph which is non-transmission regular. First of all, we will give some concepts and lemmas.
Lemma 2.1.
(Perron-Frobenius) ([10]) Let $M$ be irreducible and nonnegative square matrix of order $n$. Let $\rho(M)$ be the spectral radius of $M$. Then the following results hold.
(a) $\rho(M)>0$;
(b) $\rho(M)$ is an algebraically and geometrically simple eigenvalue of $M$;
(c) There is a positive eigenvector corresponding to eigenvalue $\rho(M)$.
Then we give the definition of the equitable quotient matrix and its some properties.
Lemma 2.2.
([22]) Let $M$ be a complex matrix of order $n$ described in the following block form
$M=\begin{bmatrix}M_{11}&\cdots&M_{1t}\\ \vdots&\ddots&\vdots\\ M_{t1}&\cdots&M_{tt}\end{bmatrix},$
where the blocks $M_{ij}$ are $n_{i}\times n_{j}$ matrices for any $1\leq i,j\leq t$ and $n=n_{1}+n_{2}+\cdots+n_{t}.$ For $1\leq i,j\leq t$, let $b_{ij}$ denote the average row sum of $M_{ij}$, i.e. $b_{ij}$ is the sum of all entries in $M_{ij}$ divided by the number of rows. Then $B(M)=(b_{ij})$(or simply $B$) is called the quotient matrix of $M$. If, in addition, for each pair $i$, $j$, $M_{ij}$ has a constant row sum, then $B$ is called the equitable quotient matrix of $M$.
Lemma 2.3.
([22]) Let $B$ be the equitable quotient matrix of $M$. In addition, let $M$ be a nonnegative matrix. Then the spectral radius relation satisfies $\rho(B)=\rho(M)$.
For a connected graph $G$ with order $n$, take $\tau(G)=\frac{Tr_{max}(G)-\mu_{\alpha}(G)}{1-\alpha}$, where $Tr_{max}$ is the maximum vertex transmission of $G$. Then the difference between the maximum transmission and the generalized distance spectral radius of $G$ is $(1-\alpha)\tau(G)$.
In [3], Atik and Panigrahi introduced a class of graphs called $DVDR$ graphs. Let $G$ be a connected graph with order $n$. If there exist a vertex $v$ in $G$ such that the degree $d(v)=n-1$ and $G-v$ is an regular graph, then we call $G$ to be a distinguished vertex deleted regular graph ($DVDR$). The vertex $v$ with $d(v)=n-1$ is said to be distinguished vertex of graph $G$. In addition, we say $G$ is an $r$-$DVDR$ graph if $G$-$v$ is $r$-$regular$.
Let $G_{0}$ be a graph which attains the minimum of $Tr_{max}(G)-\mu_{\alpha}(G)$ among all connected non-transmission regular graph $G$ of order $n$. Since $G_{0}$ is a connected graph, then it follows from the Perron Frobenius Theorem [10] that $\mu_{\alpha}(G_{0})$ is an eigenvalue of $D_{\alpha}(G_{0})$, and there is a unique positive unit eigenvector corresponding to $\mu_{\alpha}(G_{0})$. Let $X=(x_{1},x_{2},\cdots,x_{n})$ be the unique positive unit eigenvector corresponding to $\mu_{\alpha}(G_{0})$. To simplify the writing process, let $x_{max}=max\{x_{i}|1\leq i\leq n\}$ and $x_{min}=min\{{x_{i}|1\leq i\leq n\}}$. Let $Tr_{max}=max\{Tr(v)|v\in V(G_{0})\}$ and $W=W(G_{0})=\frac{1}{2}\sum\limits_{w\in V(G_{0})}Tr_{w}$. By a simple calculation, we get
$Tr_{max}-\mu_{\alpha}(G_{0})$ $=X^{T}(Tr_{max}I_{n})X-X^{T}D_{\alpha}(G_{0})X$ (1)
$=Tr_{max}-2(1-\alpha)\sum\limits_{1\leq i<j\leq n}d(v_{i},v_{j})x_{i}x_{j}-\alpha\sum\limits_{i=1}^{n}Tr_{i}x_{i}^{2}$
$\geq Tr_{max}-2(1-\alpha)\sum\limits_{1\leq i<j\leq n}d(v_{i},v_{j})x_{i}x_{j}-\alpha Tr_{max}\sum\limits_{i=1}^{n}x_{i}^{2}$
$=(1-\alpha)[Tr_{max}-2\sum\limits_{1\leq i<j\leq n}d(v_{i},v_{j})x_{i}x_{j}]$
$=(1-\alpha)[Tr_{max}\sum\limits_{i=1}^{n}x_{i}^{2}-\sum\limits_{i=1}^{n}Tr_{i}x_{i}^{2}+\sum\limits_{i=1}^{n}Tr_{i}x_{i}^{2}-2\sum\limits_{1\leq i<j\leq n}d(v_{i},v_{j})x_{i}x_{j}]$
$=(1-\alpha)[\sum\limits_{i=1}^{n}(Tr_{max}-Tr_{i})x_{i}^{2}+\sum\limits_{1\leq i<j\leq n}d(v_{i},v_{j})(x_{i}-x_{j})^{2}]$
$\geq(1-\alpha)[(nTr_{max}-2W)x_{min}^{2}+\sum\limits_{1\leq i<j\leq n}d(v_{i},v_{j})(x_{i}-x_{j})^{2}],$
where $I_{n}$ is an identity matrix of order $n$.
In the following, we will calculate the $(1-\alpha)\tau(G)$ of two special graphs.
Lemma 2.4.
Let $n=2r-1$ and $K=K_{1,\underbrace{2,2,\cdots,2}_{r-1}}$ be the complete $r$-$partite$ graph on $n$ vertices which all subsets contain two vertices except one subset contains only one vertex $v$ with $d(v)=n-1$. Then
$(1-\alpha)\tau(K)=Tr_{max}(K)-\mu_{\alpha}(K)=\frac{(1-\alpha)n+1-\sqrt{[(1-\alpha)n+1]^{2}-4(1-\alpha)}}{2}.$ (2)
Proof.
Let $v$ be the distinguished vertex of $V(K)$. Then the equitable quotient matrix of $D_{\alpha}(K)$ with respect to the partition ${\{v\}}\cup(V(K)\setminus\{v\})$ is
$M=\begin{bmatrix}\alpha(n-1)&(1-\alpha)(n-1)\\ 1-\alpha&\alpha n+(1-\alpha)(n-1)\end{bmatrix}.$
By Lemma 2.3, we know that $\mu_{\alpha}(K)$ is equal to the largest eigenvalue of $M$. Hence, we have
$\mu_{\alpha}(K)=\frac{(1+\alpha)n-1+\sqrt{[(1-\alpha)n+1]^{2}-4(1-\alpha)}}{2}.$
Since $Tr_{max}(K)=n$, by a simple calculation , we get
$(1-\alpha)\tau(K)=Tr_{max}(K)-\mu_{\alpha}(K)=\frac{(1-\alpha)n+1-\sqrt{[(1-\alpha)n+1]^{2}-4(1-\alpha)}}{2},$
which completes the proof. ∎
Lemma 2.5.
Let $G$ be an $(n-4)$-$DVDR$ graph of order $n$. Let $v$ be the distinguished vertex with $d(v)=n-1$. Then
$(1-\alpha)\tau(G)=Tr_{max}(G)-\mu_{\alpha}(G)=\frac{(1-\alpha)n+2-\sqrt{[(1-\alpha)n+2]^{2}-8(1-\alpha)}}{2}.$ (3)
Proof.
Similar to the proof of Lemma 2.4, the equitable quotient matrix of $D_{\alpha}(G)$ with respect to the partition ${\{v\}}\cup(V(G)\setminus\{v\})$ is
$N=\begin{bmatrix}\alpha(n-1)&(1-\alpha)(n-1)\\ 1-\alpha&\alpha(n+1)+(1-\alpha)n\end{bmatrix}.$
By Lemma 2.3, we know that $\mu_{\alpha}(G)$ is equal to the largest eigenvalue of $N$. Hence, we have
$\mu_{\alpha}(G)=\frac{(1+\alpha)n+\sqrt{[(1-\alpha)n+2]^{2}-8(1-\alpha)}}{2}.$
Since $Tr_{max}(G)=n+1$, by a simple calculation , we get
$(1-\alpha)\tau(G)=Tr_{max}(G)-\mu_{\alpha}(G)=\frac{(1-\alpha)n+2-\sqrt{[(1-\alpha)n+2]^{2}-8(1-\alpha)}}{2},$
completing the proof. ∎
Based on the above results, we define:
$(1-\alpha)\tau_{n}=\frac{(1-\alpha)n+\rho_{n}-\sqrt{[(1-\alpha)n+\rho_{n}]^{2}-4\rho_{n}(1-\alpha)}}{2},$ (4)
where $\rho_{n}=1$ if $n$ is odd; $\rho_{n}=2$ if $n$ is even. Clearly, $\tau_{n}$ satisfies the following equation:
$(1-\alpha)\tau_{n}^{2}-[(1-\alpha)n+\rho_{n}]\tau_{n}+\rho_{n}=0$ (5)
It is obvious that if $n$ is odd, then $(1-\alpha)\tau_{n}=(1-\alpha)\tau(K_{1,2,2,\cdots,2})$; if $n$ is even, then $(1-\alpha)\tau_{n}=(1-\alpha)\tau(G)$ for any $(n-4)$-$DVDR$ graph $G$. Noting that the above graphs are both connected non-transmission regular graphs and $G_{0}$ be a graph which attains the minimum of $(1-\alpha)\tau(G)$ among all connected non-transmission regular graph $G$ of order $n$. Hence, we have $\tau(G_{0})\leq\tau(G)$. Specifically, $\tau(G_{0})\leq\tau_{n}$.
Theorem 2.6.
Let $G$ be a connected non-transmission regular graph of order $n$. (1) If $n$ is odd and $\alpha\in[0,1)$, then
$Tr_{max}-\mu_{\alpha}(G)\geq\frac{(1-\alpha)n+1-\sqrt{[(1-\alpha)n+1]^{2}-4(1-\alpha)}}{2},$ (6)
equality holds if and only if $G\cong K_{1,2,2,\cdots,2}.$ (2) If $n$ is even and $\alpha\in[0,1)$, then
$Tr_{max}-\mu_{\alpha}(G)\geq\frac{(1-\alpha)n+2-\sqrt{[(1-\alpha)n+2]^{2}-8(1-\alpha)}}{2},$ (7)
equality holds if and only if $G$ is isomorphic to any $(n-4)$-$DVDR$ graph.
Proof.
According to the previous note on $G_{0}$, let $u$ and $v$ be two vertices of $G_{0}$ such that $x_{u}=x_{max}$ and $x_{v}=x_{min}$, respectively. Let $\mu_{\alpha}=\mu_{\alpha}(G_{0})$. From the $u$-th equation of $D_{\alpha}(G_{0})X=\mu_{\alpha}X$, we have
$\mu_{\alpha}x_{u}$ $=\alpha Tr_{u}x_{u}+(1-\alpha)\sum_{w\in V(G_{0})}d(u,w)x_{w}$
$=\alpha Tr_{u}x_{u}+(1-\alpha)[\sum_{w\in V(G_{0})-\{v\}}d(u,w)x_{w}+d(u,v)x_{v}]$
$\leq\alpha Tr_{max}x_{u}+(1-\alpha)[(Tr_{max}-d(u,v))x_{u}+d(u,v)x_{v}]$
$=Tr_{max}x_{u}+(1-\alpha)[-d(u,v)x_{u}+d(u,v)x_{v}].$
Based on the above result, we have
$Tr_{max}-\mu_{\alpha}\geq(1-\alpha)d(u,v)(1-\frac{x_{v}}{x_{u}})\geq(1-\alpha)(1-\frac{x_{v}}{x_{u}}).$
In addition, according to the properties of $G_{0}$, $\tau(G_{0})=\frac{Tr_{\max}-\mu_{\alpha}}{1-\alpha}\leq\tau_{n}$. Hence, we obtain
$(1-\tau_{n})x_{u}\leq x_{v}.$
Claim 1: $\tau_{n}<1$.
Proof: From (4), we consider two cases:
(1) If $n$ is odd, then $\rho_{n}$=1. Thus, $\tau_{n}=\frac{(1-\alpha)n+1-\sqrt{[(1-\alpha)n+1]^{2}-4(1-\alpha)}}{2(1-\alpha)}$. We take $\tau_{n}(1)=\tau_{n}$ under these conditions. If $\tau_{n}(1)\geq 1$, by a simple calculation, we get $(1-\alpha)^{2}(n-2)^{2}\geq(1-\alpha)^{2}n^{2}$, this is a contradiction.
(2) If $n$ is even, then $\rho_{n}$=2. Thus, $\tau_{n}=\frac{(1-\alpha)n+2-\sqrt{[(1-\alpha)n+2]^{2}-8(1-\alpha)}}{2(1-\alpha)}$. We take $\tau_{n}(2)=\tau_{n}$ under these conditions. Using the same method as (1), we get $\tau_{n}(2)<1$.
Hence, $\tau_{n}<1$, we get
$\frac{x_{max}}{x_{min}}=\frac{x_{u}}{x_{v}}\leq\frac{1}{1-\tau_{n}}.$ (8)
Claim 2: $nTr_{max}-2W=\rho_{n}$.
Proof: Note that $D_{\alpha}(G_{0})X=\mu_{\alpha}X$, we have $\mu_{\alpha}(1,1,\cdots,1)X=(1,1,\cdots,1)D_{\alpha}(G)X$. Furthermore,
$\mu_{\alpha}\sum\limits_{w\in V(G_{0})}x_{w}=\alpha\sum\limits_{w\in V(G_{0})}Tr_{w}x_{w}+(1-\alpha)\sum\limits_{w\in V(G_{0})}Tr_{w}x_{w}=\sum\limits_{w\in V(G_{0})}Tr_{w}x_{w}.$
Thus, we have
$(Tr_{max}-\mu_{\alpha})\sum_{w\in V(G_{0})}x_{w}=\sum_{w\in V(G_{0})}(Tr_{max}-Tr_{w})x_{w}\\ \geq(nTr_{max}-2W)x_{min}.$ (9)
Therefore, we get that
$nTr_{max}-2W\leq\frac{(Tr_{max}-\mu_{\alpha})\sum\limits_{w\in V(G_{0})}x_{w}}{x_{min}}.\$ (10)
By (8), since $Tr_{max}-\mu_{\alpha}=(1-\alpha)\tau(G_{0})\leq(1-\alpha)\tau_{n}$ and $\sum\limits_{w\in V(G_{0})}x_{w}\leq x_{min}+(n-1)x_{max}$, we have that
$nTr_{max}-2W$ $\leq\frac{(Tr_{max}-\mu_{\alpha})\sum_{w\in V(G_{0})}x_{w}}{x_{min}}$
$\leq\frac{(1-\alpha)\tau_{n}[x_{min}+(n-1)x_{max}]}{x_{min}}$
$=(1-\alpha)\tau_{n}[1+(n-1)\frac{x_{max}}{x_{min}}]$
$\leq(1-\alpha)\tau_{n}[1+(n-1)\frac{1}{1-\tau_{n}}]$
$=(1-\alpha)\tau_{n}\frac{n-\tau_{n}}{1-\tau_{n}}.$
Through (5), we have
$\rho_{n}=\frac{(1-\alpha)(n\tau_{n}-\tau^{2}_{n})}{1-\tau_{n}}.$ (11)
Thus we have
$nTr_{max}-2W\leq\rho_{n}.$
Then we will check that $nTr_{max}-2W\geq\rho_{n}$. Note that $nTr_{max}-2W$ is a positive integer. If $n$ is odd, we have $\rho_{n}=1$, it is obvious that $nTr_{max}-2W\geq 1=\rho_{n}$. If $n$ is even, we have $\rho_{n}=2$ and $nTr_{max}-2W$ is an even positive integer. Thus we have $nTr_{max}-2W\geq 2=\rho_{n}$. Thus, we have $nTr_{max}-2W=\rho_{n}$. Claim 3: $Tr_{v}<Tr_{max}$ and $Tr_{u}>Tr_{min}$.
Proof: Since $D_{0}(G)X=\mu_{\alpha}X$, from the $v$-th equation, we get
$\mu_{\alpha}x_{v}=\alpha Tr_{v}x_{v}+(1-\alpha)\sum\limits_{w\in V(G_{0})}d(v,w)x_{w}\geq Tr_{v}x_{v},$
hence $Tr_{v}\leq\mu_{\alpha}$. Since $\mu_{\alpha}<Tr_{max}$, thus $Tr_{v}<Tr_{max}$. Similarly, we get that $Tr_{u}>Tr_{min}$. Claim 4: $Tr_{max}-\mu_{\alpha}\geq\frac{\rho_{n}}{1+(n-1)\frac{x_{max}}{x_{min}}}.$ Proof: It follows from (9) and Claim 2, we have that
$(Tr_{max}-\mu_{\alpha})\sum_{w\in V(G_{0})}x_{w}=\sum_{w\in V(G_{0})}(Tr_{max}-Tr_{w})x_{w}\\ \geq(nTr_{max}-2W)x_{min}=\rho_{n}x_{min}.$ (12)
Hence, we have
$Tr_{max}-\mu_{\alpha}\geq\frac{\rho_{n}x_{min}}{\sum\limits_{w\in V(G_{0})}x_{w}}\geq\frac{\rho_{n}x_{min}}{x_{min}+(n-1)x_{max}}=\frac{\rho_{n}}{1+(n-1)\frac{x_{max}}{x_{min}}}.$ (13)
Since $\frac{x_{max}}{x_{min}}\leq\frac{1}{1-\tau_{n}}$, we have
$Tr_{max}-\mu_{\alpha}\geq\frac{\rho_{n}}{1+(n-1)\frac{1}{1-\tau_{n}}}=\frac{\rho_{n}(1-\tau_{n})}{n-\tau_{n}}.$ (14)
In addition, from (5), we get
$\rho_{n}(\tau_{n}-1)=(1-\alpha)(\tau^{2}_{n}-n\tau_{n}),$
hence (14) can be written simply as
$Tr_{max}-\mu_{\alpha}\geq\frac{\rho_{n}(1-\tau_{n})}{n-\tau_{n}}=(1-\alpha)\tau_{n}.$ (15)
Claim 5: If the equality in (14) holds, then $Tr_{min}=n-1$ and $Tr_{max}=Tr_{min}+\rho_{n}=n-1+\rho_{n}.$
Proof: If the equality in (14) holds, then the above equations hold. Furthermore, we get that
$\frac{x_{u}}{x_{v}}=\frac{x_{max}}{x_{min}}=\frac{1}{1-\tau_{n}}.$ (16)
From (13), we have $x_{w}=x_{max}$ for $w\in V(G_{0})-\{v\}$. Since $x_{v}=x_{min}$, from (12), we have
$\sum_{w\in V(G_{0})}(Tr_{max}-Tr_{w})x_{w}=\sum_{w\in V(G_{0})-\{v\}}(Tr_{max}-Tr_{w})x_{max}+(Tr_{max}-Tr_{v})x_{v}=\rho_{n}x_{min}.$
Then we will proof that $Tr_{w}=Tr_{max}$ for ${w\in V(G_{0})-\{v\}}$. If it does not true, we have $\sum\limits_{w\in V(G_{0})-\{v\}}(Tr_{max}-Tr_{w})x_{max}\geq x_{max}$. From Claim 3, we get $(Tr_{max}-Tr_{v})x_{v}\geq x_{v}=x_{min}.$ Since $G_{0}$ is non-transmission regular, then we have $x_{max}>x_{min}$. Hence
$\sum_{w\in V(G_{0})-\{v\}}(Tr_{max}-Tr_{w})x_{max}+(Tr_{max}-Tr_{v})x_{v}\geq x_{max}+x_{min}>2x_{min}\geq\rho_{n}x_{min}.$
This is a contradiction. Hence, we get that $Tr_{w}=Tr_{max}$ for ${w\in V(G_{0})-\{v\}}$ and
$Tr_{max}-Tr_{v}=Tr_{max}-Tr_{min}=\rho_{n}.$ (17)
Hence
$\mu_{\alpha}x_{v}$ $=\alpha Tr_{v}x_{v}+(1-\alpha)\sum\limits_{w\in V(G_{0})}d(v,w)x_{w}$
$=\alpha Tr_{min}x_{v}+(1-\alpha)Tr_{min}x_{max}.$
From (16), we get
$\mu_{\alpha}=\alpha Tr_{min}+(1-\alpha)Tr_{min}\frac{x_{u}}{x_{v}}=\alpha Tr_{min}+(1-\alpha)Tr_{min}\frac{1}{1-\tau_{n}}.$ (18)
From (14), (17) and (18), we get
$\rho_{n}\frac{1-\tau_{n}}{n-\tau_{n}}$ $=Tr_{max}-\mu_{\alpha}$
$=Tr_{max}-\alpha Tr_{min}-(1-\alpha)Tr_{min}\frac{1}{1-\tau_{n}}$
$=\rho_{n}+(1-\alpha)Tr_{min}-(1-\alpha)Tr_{min}\frac{1}{1-\tau_{n}},$
which implies $Tr_{min}=\frac{1}{1-\alpha}\frac{n-1}{n-\tau_{n}}\frac{1-\tau_{n}}{\tau_{n}}\rho_{n}.$ Since (5), we get $\rho_{n}=\frac{(1-\alpha)\tau^{2}_{n}-(1-\alpha)n\tau_{n}}{\tau_{n}-1}.$ Then
$\begin{cases}Tr_{min}=\frac{1}{1-\alpha}\frac{n-1}{n-\tau_{n}}\frac{1-\tau_{n}}{\tau_{n}}\rho_{n}\\ \rho_{n}=\frac{(1-\alpha)\tau^{2}_{n}-(1-\alpha)n\tau_{n}}{\tau_{n}-1}\end{cases},$
by a simple calculation, we get $Tr_{min}=n-1$ and $Tr_{max}=Tr_{min}+\rho_{n}=n-1+\rho_{n}$, completing the proof of this claim.
Hence, we will characterize the necessary and sufficient conditions for the equality in (14).
On the one hand, from Claim 5, we get $d(v)=n-1$ and $Tr_{min}=Tr_{v}=n-1$, which implies $dim(G)=2$. For any $y\in V(G_{0})$,
$Tr_{y}=\sum\limits_{x\in V(G_{0})}d(y,x)=d(y)+2(n-1-d(y))=2n-2-d(y).$
Hence, for $z\in V(G_{0})-\{v\}$, we have $d(z)=2(n-1)-Tr_{z}=2n-2-(n-1+\rho_{n})=n-1-\rho_{n}$, which implies that if $n$ is odd, $d(z)=n-2$; if n is even, $d(z)=n-3.$ Hence, if $n$ is odd, then $G_{0}\cong K_{1,2,\cdots,2}$; if $n$ is even, then $G_{0}$ is isomorphic to any $(n-4)$-$DVDR$ graph.
On the other hand, under the following conditions:
(1) If $n$ is odd, $\alpha\in[0,1)$ and $G\cong K_{1,2,\cdots,2}.$
(2) If $n$ is even, $\alpha\in[0,1)$ and $G_{0}$ is isomorphic to any $(n-4)$-$DVDR$ graph.
It is easy to see that
$Tr_{max}(G)-\mu_{\alpha}(G)=\frac{\rho_{n}(1-\tau_{n})}{n-\tau_{n}}=(1-\alpha)\tau_{n},$
where $\tau_{n}=\tau_{n}(1)$ if $n$ is odd and $\tau_{n}=\tau_{n}(2)$ if $n$ is even. Therefore, the equation clearly holds.
Since $G_{0}$ is a graph which attains the minimum of $Tr_{max}(G)-\mu_{\alpha}(G)$ among all connected non-transmission regular graph $G$ of order $n$. Hence, for any connected non-transmission regular graph $G$, we have:
$Tr_{max}(G)-\mu_{\alpha}(G)\geq Tr_{max}(G_{0})-\mu_{\alpha}(G_{0})\geq(1-\alpha)\tau_{n}.$
According to the above process, we have:
(1) If $n$ is odd and $\alpha\in[0,1)$, then
$Tr_{max}-\mu_{\alpha}(G)\geq(1-\alpha)\tau_{n}(1)=\frac{(1-\alpha)n+1-\sqrt{[(1-\alpha)n+1]^{2}-4(1-\alpha)}}{2},$
where $\tau_{n}(1)=\frac{(1-\alpha)n+1-\sqrt{[(1-\alpha)n+1]^{2}-4(1-\alpha)}}{2(1-\alpha)}.$ Equality holds if and only if $G\cong K_{1,2,\cdots,2}.$ (2) If $n$ is even and $\alpha\in[0,1)$, then
$Tr_{max}-\mu_{\alpha}(G)\geq(1-\tau_{n}(2))=\tau_{n}(2)=\frac{(1-\alpha)n+2-\sqrt{[(1-\alpha)n+2]^{2}-8(1-\alpha)}}{2},$
where $\tau_{n}(2)=\frac{(1-\alpha)n+2-\sqrt{[(1-\alpha)n+2]^{2}-8(1-\alpha)}}{2(1-\alpha)}.$ Equality holds if and only if $G$ is isomorphic to any $(n-4)$-$DVDR$ graph. ∎
In the following, let $\alpha=0$, then $\tau_{n}(1)=\frac{n+1-\sqrt{(n+1)^{2}-4}}{2}$ and $\tau_{n}(2)=\frac{n+2-\sqrt{(n+2)^{2}-8}}{2}.$ Hence, we can get the same conclusion as Theorem 1.1 in [15].
Corollary 2.7.
([15]) Let $G$ be a connected non-transmission regular graph of order $n$. Let $\mu(G)$ be the distance spectral radius of $G$, then (1) If $n$ is odd, then
$Tr_{max}-\mu(G)\geq\frac{n+1-\sqrt{(n+1)^{2}-4}}{2}.$ (19)
Equality holds if and only if $G\cong K_{1,2,\cdots,2}.$ (2) If $n$ is even, then
$Tr_{max}-\mu(G)\geq\frac{n+2-\sqrt{(n+2)^{2}-8}}{2}.$ (20)
Equality holds if and only if $G$ is isomorphic to any $(n-4)$-$DVDR$ graph.
In the following corollary, let $\alpha=\frac{1}{2}$, then $\tau_{n}(1)=\frac{n+2-\sqrt{n^{2}+4n-4}}{2}$ and $\tau_{n}(2)=\frac{n+4-\sqrt{n^{2}+8n}}{2}.$ Hence, we can get the same conclusion as Theorem 1.1 in [11].
Corollary 2.8.
([15]) Let $G$ be a connected non-transmission regular graph of order $n$. Let $\mu_{Q}(G)$ be the distance signless Laplacian spectral radius of $G$, then (1) If $n$ is odd, then
$2Tr_{max}-\mu_{Q}(G)\geq\frac{n+2-\sqrt{n^{2}+4n-4}}{2}.$ (21)
Equality holds if and only if $G\cong K_{1,2,\cdots,2}.$ (2) If $n$ is even, then
$2Tr_{max}-\mu_{Q}(G)\geq\frac{n+4-\sqrt{n^{2}+8n}}{2}.$ (22)
Equality holds if and only if $G$ is isomorphic to any $(n-4)$-$DVDR$ graph.